Please help me solve this problem:
A 2-digit number is twice the difference of its digits. What is the number?

Answer by Edwin McCravy(20056)

Let t = the tens digit
Let u = the units digit.

Then the number = 10t+u
The difference of digits is t-u
Twice the difference of digits = 2(t-u)

We want cases of

10t+u = 2(t-u)

10t+u = 2t-2u

8t+3u = 0

That certainly is not possible.

Perhaps you mean to subtract them the other way u-t

10t+u = 2(u-t)

10t+u = 2u-2t

  12t = u 

That's certainly not possible, either
 
Here is a computer-generated list to show that there is no 
such two-digit number:

Twice the difference of digits of 10 is 2(1-0) = 2 not 10
Twice the difference of digits of 11 is 2(1-1) = 0 not 11
Twice the difference of digits of 12 is 2(1-2) = -2 not 12
Twice the difference of digits of 13 is 2(1-3) = -4 not 13
Twice the difference of digits of 14 is 2(1-4) = -6 not 14
Twice the difference of digits of 15 is 2(1-5) = -8 not 15
Twice the difference of digits of 16 is 2(1-6) = -10 not 16
Twice the difference of digits of 17 is 2(1-7) = -12 not 17
Twice the difference of digits of 18 is 2(1-8) = -14 not 18
Twice the difference of digits of 19 is 2(1-9) = -16 not 19
Twice the difference of digits of 20 is 2(2-0) = 4 not 20
Twice the difference of digits of 21 is 2(2-1) = 2 not 21
Twice the difference of digits of 22 is 2(2-2) = 0 not 22
Twice the difference of digits of 23 is 2(2-3) = -2 not 23
Twice the difference of digits of 24 is 2(2-4) = -4 not 24
Twice the difference of digits of 25 is 2(2-5) = -6 not 25
Twice the difference of digits of 26 is 2(2-6) = -8 not 26
Twice the difference of digits of 27 is 2(2-7) = -10 not 27
Twice the difference of digits of 28 is 2(2-8) = -12 not 28
Twice the difference of digits of 29 is 2(2-9) = -14 not 29
Twice the difference of digits of 30 is 2(3-0) = 6 not 30
Twice the difference of digits of 31 is 2(3-1) = 4 not 31
Twice the difference of digits of 32 is 2(3-2) = 2 not 32
Twice the difference of digits of 33 is 2(3-3) = 0 not 33
Twice the difference of digits of 34 is 2(3-4) = -2 not 34
Twice the difference of digits of 35 is 2(3-5) = -4 not 35
Twice the difference of digits of 36 is 2(3-6) = -6 not 36
Twice the difference of digits of 37 is 2(3-7) = -8 not 37
Twice the difference of digits of 38 is 2(3-8) = -10 not 38
Twice the difference of digits of 39 is 2(3-9) = -12 not 39
Twice the difference of digits of 40 is 2(4-0) = 8 not 40
Twice the difference of digits of 41 is 2(4-1) = 6 not 41
Twice the difference of digits of 42 is 2(4-2) = 4 not 42
Twice the difference of digits of 43 is 2(4-3) = 2 not 43
Twice the difference of digits of 44 is 2(4-4) = 0 not 44
Twice the difference of digits of 45 is 2(4-5) = -2 not 45
Twice the difference of digits of 46 is 2(4-6) = -4 not 46
Twice the difference of digits of 47 is 2(4-7) = -6 not 47
Twice the difference of digits of 48 is 2(4-8) = -8 not 48
Twice the difference of digits of 49 is 2(4-9) = -10 not 49
Twice the difference of digits of 50 is 2(5-0) = 10 not 50
Twice the difference of digits of 51 is 2(5-1) = 8 not 51
Twice the difference of digits of 52 is 2(5-2) = 6 not 52
Twice the difference of digits of 53 is 2(5-3) = 4 not 53
Twice the difference of digits of 54 is 2(5-4) = 2 not 54
Twice the difference of digits of 55 is 2(5-5) = 0 not 55
Twice the difference of digits of 56 is 2(5-6) = -2 not 56
Twice the difference of digits of 57 is 2(5-7) = -4 not 57
Twice the difference of digits of 58 is 2(5-8) = -6 not 58
Twice the difference of digits of 59 is 2(5-9) = -8 not 59
Twice the difference of digits of 60 is 2(6-0) = 12 not 60
Twice the difference of digits of 61 is 2(6-1) = 10 not 61
Twice the difference of digits of 62 is 2(6-2) = 8 not 62
Twice the difference of digits of 63 is 2(6-3) = 6 not 63
Twice the difference of digits of 64 is 2(6-4) = 4 not 64
Twice the difference of digits of 65 is 2(6-5) = 2 not 65
Twice the difference of digits of 66 is 2(6-6) = 0 not 66
Twice the difference of digits of 67 is 2(6-7) = -2 not 67
Twice the difference of digits of 68 is 2(6-8) = -4 not 68
Twice the difference of digits of 69 is 2(6-9) = -6 not 69
Twice the difference of digits of 70 is 2(7-0) = 14 not 70
Twice the difference of digits of 71 is 2(7-1) = 12 not 71
Twice the difference of digits of 72 is 2(7-2) = 10 not 72
Twice the difference of digits of 73 is 2(7-3) = 8 not 73
Twice the difference of digits of 74 is 2(7-4) = 6 not 74
Twice the difference of digits of 75 is 2(7-5) = 4 not 75
Twice the difference of digits of 76 is 2(7-6) = 2 not 76
Twice the difference of digits of 77 is 2(7-7) = 0 not 77
Twice the difference of digits of 78 is 2(7-8) = -2 not 78
Twice the difference of digits of 79 is 2(7-9) = -4 not 79
Twice the difference of digits of 80 is 2(8-0) = 16 not 80
Twice the difference of digits of 81 is 2(8-1) = 14 not 81
Twice the difference of digits of 82 is 2(8-2) = 12 not 82
Twice the difference of digits of 83 is 2(8-3) = 10 not 83
Twice the difference of digits of 84 is 2(8-4) = 8 not 84
Twice the difference of digits of 85 is 2(8-5) = 6 not 85
Twice the difference of digits of 86 is 2(8-6) = 4 not 86
Twice the difference of digits of 87 is 2(8-7) = 2 not 87
Twice the difference of digits of 88 is 2(8-8) = 0 not 88
Twice the difference of digits of 89 is 2(8-9) = -2 not 89
Twice the difference of digits of 90 is 2(9-0) = 18 not 90
Twice the difference of digits of 91 is 2(9-1) = 16 not 91
Twice the difference of digits of 92 is 2(9-2) = 14 not 92
Twice the difference of digits of 93 is 2(9-3) = 12 not 93
Twice the difference of digits of 94 is 2(9-4) = 10 not 94
Twice the difference of digits of 95 is 2(9-5) = 8 not 95
Twice the difference of digits of 96 is 2(9-6) = 6 not 96
Twice the difference of digits of 97 is 2(9-7) = 4 not 97
Twice the difference of digits of 98 is 2(9-8) = 2 not 98
Twice the difference of digits of 99 is 2(9-9) = 0 not 99

Edwin